OpenStudy (anonymous):

Sam has a total of 58 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have? 42 14 44 12

OpenStudy (chrisdbest):

14 is my lucky guess

OpenStudy (anonymous):

let the no. of dvds be x. by problem, (2+3x)+x=58 4x=56 x=56/4 x=14. no. of dvds=14 no. of cds=58-14=44

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

have u undrstood?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

ok then...wlcome!!

OpenStudy (chrisdbest):

did you get the q wrong?

OpenStudy (anonymous):

idk havnt submitted it yet lol have 3 other questions left

OpenStudy (chrisdbest):

hit me with them

OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the rate in feet per second.

OpenStudy (anonymous):

final speed is zero.

OpenStudy (anonymous):

i need help solving it though :/

OpenStudy (anonymous):

The height in meters of a projectile can be modeled by h = −4.9t2 + vt + s where t is the time (in seconds) the object has been in the air, v is the initial velocity (in meters per seconds) and s is the initial height (in meters). A soccer ball is kicked upward from the ground and flies through the air with an initial vertical velocity of 4.9 meters per second. After how many seconds does it land approximately?

OpenStudy (chrisdbest):

is that baseball one is 16+2 or 16*2?

OpenStudy (anonymous):

what? like the equation?

OpenStudy (chrisdbest):

yes

OpenStudy (anonymous):

16x^2

OpenStudy (anonymous):

so the x is squared

OpenStudy (anonymous):

@chrisdbest

OpenStudy (chrisdbest):

ok